# POV Learning: Individual Alignment of Multimodal Models using Human Perception

### v1

@@ Abstract
Aligning machine learning systems with human expectations is mostly attempted by training with manually vetted human behavioral samples, typically explicit feedback. This is done on a population level since the context that is capturing the subjective Point-Of-View (POV) of a concrete person in a specific situational context is not retained in the data.
However, we argue that alignment on an individual level can boost the subjective predictive performance for the individual user interacting with the system considerably. Since perception differs for each person, the same situation is observed differently. Consequently, the basis for decision making and the subsequent reasoning processes and observable reactions differ.
We hypothesize that individual perception patterns can be used for improving the alignment on an individual level. We test this, by integrating perception information into machine learning systems and measuring their predictive performance wrt. individual subjective assessments. For our empirical study, we collect a novel data set of multimodal stimuli and corresponding eye tracking sequences for the novel task of Perception-Guided Crossmodal Entailment and tackle it with our Perception-Guided Multimodal Transformer.
Our findings suggest that exploiting individual perception signals for the machine learning of subjective human assessments provides a valuable cue for individual alignment. It does not only improve the overall predictive performance from the point-of-view of the individual user but might also contribute to steering AI systems towards every person’s individual expectations and values.
@@ 1 Motivation
Since the release of GPT4 [15] multimodal transformers have received a great amount of public attention.
They show impressive capabilities on tasks like visual question answering that take image and text as input and generated text as output. However, aligning machine learning models with an individual person’s point-of-view (POV) and expectations has not attracted much research yet.
To improve the alignment of machine learning models according to individual expectations and preferences we first have to capture personality cues that are potential indicators for personal preferences and values. For a stimulus that consist of text and images the visual perception process is a potential signal that indicates an individual’s evaluation of said stimulus (see Fig. 1).
Figure 1: Visualisation of human visual perception when given the task: Are the central objects in the image mentioned in the caption? The red areas are showing the focus of human attention.
The other required component is a machine learning objective that is an indicator for individual alignment. When accounting for tasks that measure individual behavior, a key feature is that the same stimuli might be differently assessed by different individuals. Contrary to traditional machine learning tasks, this is not regarded as noise or an undesirable low inter-rater reliability in POV learning. Instead we are aiming for individual alignment, so individually differing assessments of the same stimuli is a key feature to be exploited.
In summary, a POV machine learning task as illustrated in Fig. 2 is about maximizing the POV-specific objective (here, crossmodel entailment), given the stimulus (here, text+image pairs) plus the individual alignment signal (here, eye-tracking perception traces).
@@
    For measuring individual alignment, we propose the novel POV-task of Perception-guided Crossmodal Entailment (PCE) which attempts to predict how a test person assesses the coherence of different modalities of a multimodal stimulus (e.g., an image with captions) given the test person’s visual fixation sequence. We collect the first POV learning benchmark data set based on PCE by tracing each test person’s eye movements while this person is judging if the subjectively central visual objects are mentioned in the image’s captions (see Sec. 3).
@@
    We empirically show how eye tracking sequences can be used for improved individual alignment of POV machine learning models. For this, we explore multiple ways of exploiting the eye tracking data in a variety of machine learning models (see Sec. 4 and 5). To this end, we propose the Perception-Guided Multimodal Transformer (PGMT) which is optimized for PCE (see Sec. 6).
@@
    Additionally, we assess how multimodal foundation models (i.e., GPT4) perform on POV-learning in zero and few-shot settings (see Sec. 7).
Our findings suggest that augmenting machine learning models with individual perception signals and optimizing for POV-specific objectives is a promising direction of future research. Our PGMT clearly outperforms GPT4 on the PCE task and contributes to the alignment of AI systems on an individual level. We call this new perspective on AI alignment POV Learning.
@@ 2 Related work
A variety of research areas are related to the work we present. On a technical level, the field of multi-modal (in our case visual-linguistic) representation learning provides the technical background for our experiments. The novel PCE task that is introduced in this work is related to Visual Entailment tasks and adds a behavioral component to it. Other related work will be covered in the following subsection, where we discuss previous work on eye tracking in psychology and how it relates to POV learning.
2.1 Foundation Models and their Limitations
There are two key techniques on which many recent language models are built on: the attention mechanism [1] and the transformer architecture [24]. Despite their impressive capabilities, there are downsides that come along with models like Gemini [23], GPT4 [15], Meta Llama 3111https://ai.meta.com/blog/meta-llama-3/ and beyond. Firstly, the amount of resources needed to train a modern language model in both time and storage is tremendous. Secondly, the alignment of Large Language Models (LLMs) on an individual level as a POV learning task is an open research question.
@@ 2.2 Visual-Linguistic Transformers and Tasks
There exists a variety of recent approaches in Visual-linguistic representation learning. Before the introduction of novel multi modal LLMs like GPT4 [15] and Gemini [23], which are currently the dominant approaches and outperform the traditional models across all benchmarks, they usually combined the BERT [4] architecture with a vision module. Models that fall in this category are VilBert [14], VL-Bert [20] and Visual Bert [11].
Others, like LXMERT [21], Uniter [3], SimvLM [26] and, most recently, OFA [25] use other transformer-based architectures. They all have in common that they are evaluated at least partly on one or more of the benchmark tasks provided by datasets like Visual Genome [10], MS-Coco [12], or Flickr30k [19]. Normally, the tasks are divided into two sub-categories which are briefly described here.
@@
This list is not exhaustive, but contains tasks most commonly referred to in Visual-Linguistic representation learning. The task with the highest similarity to the Perception-guided Crossmodal Entailment task we propose in this work is Visual Entailment [28]. In Visual Entailment, an image text combination functions as a premise/hypothesis pair and the task is to classify whether the hypothesis is entailed in the premise, contradicts the premise, or is neutral with regard to the premise.
@@ 2.3 AI Alignment
The goal of AI Alignment is to ensure that AI models behave in a way that overlaps with human values. Due to the ever increasing importance of LLMs and AI models in general, behavior like untruthful responses, hallucinating, or societal biases needs to be mitigated to prevent harmful effects on society [8]. To prevent such harmful responses, originally reinforcement learning from human feedback (RLHF) was introduced in the deployment process of LLMs. This technique relies on human provided feedback, which serves as an alignment signal to prevent LLMs from returning harmful output [16] and has since seen continuous improvements.
@@ 2.4 Machine Learning on Eye Tracking Data
Historically, eye tracking data has been most common as a signal in the psychology of (visual) perception, where it has been used to determine how visual perception works, both physiologically and psychologically [5]. In other lines of psychology related work, eye tracking data is used to to train statistical models for predicting behavior in binary choice settings (cmp. [9] [22] [2]). In contrast to our work, the works described above use directly derived information from the eye tracking sequences (i.e. dwell time, revisits) [18] as input for their statistical models, while we focus on symbolic representations of multi-modal AOIs [^2] In our case, we describe all annotated regions of the stimuli as AOI. These can be both the words in the caption and the visual entities in the image. that we use as input for deep learning models.
@@
Most importantly, no previous work attempts to fuse the two research areas in order to derive individually aligned machine learning models.
@@ 3 Perception-guided Crossmodal Entailment
This section describes how the data for both, the perception-guided ML model based on eye tracking data, as well as the content based transformer models were obtained. For both models the identical prediction targets are used, namely the task of Perception-guided Crossmodal Entailment (PCE).
As basis for the eye tracking study, we selected a set of multi-modal documents from the Visual Genome data set [10]. Visual Genome (VG) consists of 108,249 images that are annotated on several layers, from regions to objects, represented in a graph structures. We leverage the underlying scene graph to rank entities depicted in an image by their centrality degree. Based on this ranking, we chose regions that are connected by their contained entities and retrieve the associated caption. We extract up to three regions for each picture which, due to the procedure just described, are connected. The region captions then serve as textual description of the picture. The region descriptions combined with the image constitute the multi-modal document used for further annotation.
@@ 3.1 POV Task Specification
To incentivize the test persons we showed the question “Are the central objects in the image mentioned in the caption?” to the study’s participants since it requires crossmodal reasoning and it assesses human perception.
For the eye tracking study 153 stimuli were generated. Errors in the captions and object annotations were corrected. Objects with incorrect VG-annotations were removed. This was documented so that all preprocessing steps can be reconstructed afterwards.
@@ 3.2 Eye tracking and Indvidual Assessment Recording
Overall, 109 participants in the age range of 19-61 years participated in the study, which took place in the eye tracking laboratory of the Media Studies department at our institution. The majority of the participants were female (F= 77, 70.64 %; M= 30, 27.52 %), the average age was Mage = 25.4 years. They were recruited via several open calls on university newsletter sites, via online advertisements on social media (Facebook, Instagram), and on printed posters on campus. The experiments were conducted between 5 May and 14 July, 2022. A small participant remuneration was offered (5 € per participant). Participants were mainly students and staff of the University. Participation requirements included basic knowledge of English to ensure they understood the textual stimuli in the experimental task.
A total of 153 text-image-stimuli were selected to be shown to the participants. To ensure that the length of an eye tracking session was not too long to retain the participant’s attention, the participants were divided into three groups: group A (Nparticipants=37; Nstimuli=50), group B (Nparticipants=35; Nstimuli=50) and group C (Nparticipants=37; Nstimuli=53). With exception of the different stimuli, all other experiment settings were the same: a short instruction slide was shown followed by 50-53 blocks consisting of a text-image-stimulus and a following question slide. The blocks of stimulus and question slides were shown in a randomized order. For the study, the experiment was programmed and analyzed in the software iMotions (Version 9.1.5). The total duration of the experimental task was approximately eight minutes (excluding the time for initial calibration).
@@ 3.3 Symbolic Fixation Sequence Extraction
From the eye tracking study we obtain the participant-specific fixation sequences for all stimuli. In addition to the information about the participant and stimulus, each fixation contains information about its location (x and y coordinates), its duration, dispersion and the associated annotated region. For instance, Table 1 lists how the sequence in Figure 3 is represented. All fixations are ordered by the time of their occurrence, meaning that they are ordered chronologically.
@@
Note that in this raw data export from iMotions, there are artifacts like in row numbers 2 and 3, where the same fixation is assigned to two different AOIs. In cases like this, we removed the non-specific entry for AOI off, which normally is used to describe a fixation that is on none of the annotated AOIs. The coordinates represent fixation centers calculated by iMotions, which is why they are floating point numbers instead of integers, as one would expect for pixel values.
Table 1: Fixation sequence for Participant EWCX and Stimulus ID 2412873.
Nr. AOI Coordinates Duration
1 vis_wall 865.44, 346.46 100.00
2 txt_wall 709.13, 29.92 158.32
3 off 709.13, 29.92 158.32
4 txt_Rock 658.71, 39.78 191.67
5 txt_wall 760.30, 32.51 258.33
18 vis_ground 976.25, 859.95 166.66
19 vis_zebra 718.85, 823.35 166.66
20 vis_wall 988.11, 431.43 250.00
@@ 3.4 Data summary
In summary, the data we obtain consist of the created stimuli (images, annotated AOIs and corresponding captions) and a participant specific fixation sequence for each stimulus. Overall, the acquired data set contains 5,400 unique fixation sequences, where each sequence represents a perception pattern of a participant generated when parsing the stimulus. The fixation sequences contain a total of 148,100 identified fixations, on average 27.42 fixations per stimulus exposure. The distribution of the prediction targets is heavily skewed, with 67.4 % of the samples belonging to the Yes class, 20.1 % to the No class, and 12.5 % to the Unclear class. The data set (including raw fixation data, stimuli, caption texts and annotated regions) and the code for the experiments is available on GitHub. [^3] https://github.com/siwer/Crossmodal-Entailment
@@ 4 POV Models and Experimental Design
To test our hypothesis that the evaluation process of stimuli is reflected in the eye tracking data and can act as a feature for improving individual alignment, three different models with focus on different aspects of our data were implemented.
The focus of the LSTM [7] model is on exploring the informational value of the fixation sequence of symbolic components of the multimodal stimulus as alignment signal, without relying on content information about the visual and textual entities in the stimuli.
In contrast to this, the focus of the Transformer model is on content-based information in form of pre-trained feature representations of the stimuli. To enable the Transformer to being able to take into account individual preferences, we also provided a rudimentary participant representation as an alignment signal.
Finally, as a combination of both, we train an Ensemble model that evaluates the predictions of both models jointly.
The task for all models is to predict each participant’s evaluation of the text-image combinations individually. This problem can be formulated as a classification problem with three possible outcomes: 1) yes, the caption mentions the central objects in the image, 2) no, central objects in the image are not mentioned and 3) unclear.
4.1 Technical Setup
The data is divided into training, evaluation and test sets in a random 80/10/10 split, with every split retaining the global class distribution from the data. All the models that are described in the following sections are optimized with AdamW [13] optimizer on a Cross-entropy loss and trained for a maximum of 30 epochs. The final model is chosen as the one with minimum loss on the evaluation split. If not stated otherwise, we use the PyTorch [17] (version 1.11) implementations for all building blocks in our models. We conduct all experiments on a GPU server at our institution.
The hyperparameter ranges that were tested to find the best performing Ensemble and best performing Perception-guided Transformer are stated in table 2. All possible combinations were evaluated. We report accuracy and F1 score as the evaluation metrics to account for the class imbalance in the data. Because of the sparsity of the unclear class, we test all models in two different settings, one with all classes present (3 classes) and one with the unclear class excluded from evaluation.
Table 2: Tested hyperparameters and their ranges.
Parameter Range
Learning Rate 1e-04, 5e-04, 1e-05, 5e-05
FF-Dim 32, 64, 128, 256
Embedding Dim 8, 16, 32
Batch size 16, 64, 128

### v2

# POV Learning: Individual Alignment of Multimodal Models via Human Perception

@@ Abstract
Aligning machine learning systems with human preferences is mostly attempted by training with manually vetted human behavioral samples, typically explicit feedback. This is done on a population level since the context that captures the subjective Point-of-View (POV) of a concrete person in a specific situational context is not retained in the data. We argue that alignment on an individual level can significantly boost subjective predictive performance for the individual user who interacts with the system. Naturally, perception differs for each person, thus the same situation is observed differently. Consequently, the basis for decision making and the subsequent reasoning processes and observable reactions differ for every individual.
We propose that individual perception patterns can be used to improve alignment on an individual level. We tested this by integrating perception information into machine learning systems and measured their predictive performance with respect to individual subjective assessments. For our empirical study, we collect a novel data set of multimodal stimuli and corresponding eye tracking sequences for the novel task of Perception-Guided Crossmodal Entailment and tackle it with our Perception-Guided Multimodal Transformer. Our findings suggest that exploiting individual perception signals for machine learning of subjective human assessments provides a valuable cue for individual alignment. It not only improves the overall predictive performance from the point of view of the individual user but might also contribute to steering AI systems towards every person’s individual preferences and values and ultimately improves human-AI interaction.
keywords:
[…]
, , , and
@@ 1 Motivation
Since the release of GPT4 [1] and open-source variants such as Pixtral [2], multimodal transformers have received a great deal of public attention. They show impressive capabilities on tasks like visual question answering that take image and text as input and generated text as output. However, aligning machine learning models with an individual person’s Point-Of-View (POV) and preferences has not attracted much research yet.
To improve the alignment of machine learning models according to individual expectations and preferences, we first have to capture personality cues that are potential indicators of personal preferences and values. For a stimulus consisting of text and images, the visual perception process is a potential signal that indicates an individual’s evaluation of the said stimulus (see Fig. 2).
[…]
Figure 1: Human visual perception when given the crossmodal entailment task: Are the central objects in the image mentioned in the caption? The red areas are indicating the focus of human attention according to eye tracking.
@@
The other required component is a machine learning objective that is an indicator of individual alignment. When accounting for tasks that measure individual behavior, a key feature is that the same stimuli might be assessed differently by different individuals. Contrary to traditional machine learning, in POV learning, this is not regarded as noise or an undesirable low inter-rater reliability. Instead, we are aiming for individual alignment, so individually differing assessments of the same stimuli is considered the key objective for optimizing the model. More formally, we define a POV machine learning task illustrated in Fig. 2 as maximizing a POV-specific objective function:
$g^{\text{\tiny POV}}(x^{\text{\tiny POV}})=\underset{y^{\text{\tiny POV}}}{\arg\max}\;f^{\text{\tiny POV}}(x^{\text{\tiny POV}},y^{\text{\tiny POV}})$
where $y^{\text{\tiny POV}}$ is a prediction target subjective to each individual person. The given features $x^{\text{\tiny POV}}$ are a tuple of the stimulus $x^{\text{\tiny STIM}}$ (here, text+image pairs) and the individual alignment signal $x^{\text{\tiny ALIGN}}$ (here, eye-tracking perception traces), thus $x^{\text{\tiny POV}}=(x^{\text{\tiny STIM}},x^{\text{\tiny ALIGN}})$.
@@
    For measuring individual alignment according to a evaluation metric $g^{\text{\tiny POV}}$, we propose the novel POV-task of Perception-guided Crossmodal Entailment (PCE) which attempts to predict how a test person assesses the coherence of different modalities of a multimodal stimulus $x^{\text{\tiny STIM}}$ (e.g., an image with captions) given the test person’s visual fixation sequence $x^{\text{\tiny ALIGN}}$. We collect the first POV learning benchmark data set based on PCE by tracing each test person’s eye movements while this person is judging if the subjectively central visual objects are mentioned in the image’s captions (see Sect. 3).
@@
    We empirically show how eye tracking sequences can be used for improved individual alignment of POV machine learning models. For this, we explore multiple ways to exploit the eye tracking data in a variety of machine learning models $f(X^{\text{\tiny POV}},Y^{\text{\tiny POV}})$ (see Sects. 4 and 5). To this end, we propose the Perception-Guided Multimodal Transformer (PGMT) which is optimized for PCE (see Sect. 6).
@@
    Furthermore, we assess how multimodal large language models (i.e., GPT4) perform on POV-learning in zero and few-shot in-context settings (see Sect. 7).
Our findings suggest that augmenting machine learning models with individual perception signals and optimizing for POV-specific objectives is a promising direction for future research. We call this new perspective on AI alignment POV Learning.
@@ 2 Related work
A variety of research areas are related to the work we present. On a technical level, the field of multi-modal (in our case visual-linguistic) representation learning provides the technical background for our experiments. The novel PCE task introduced in this work is related to the Visual Entailment tasks and adds a behavioral component to it.
2.1 Large Language Models and their Limitations
There are two key techniques on which many recent language models are based: the attention mechanism [3] and the transformer architecture [4]. Despite their impressive capabilities, there are downsides associated with models like Gemini [5], GPT4 [1], Meta Llama 3and beyond. Firstly, the amount of resources needed to train a modern language model in both time and storage is tremendous. Secondly, the alignment of Large Language Models (LLMs) on an individual level as a POV learning task is an open research question.
@@ 2.2 Visual-Linguistic Transformers and Tasks
There exists a variety of recent approaches in Visual-linguistic representation learning. Before the introduction of novel multimodal LLMs like GPT4 [1] and Gemini [5], which are currently the dominant approaches and outperform traditional models across all benchmarks, they usually combined the BERT architecture [6] with a vision module. Models that fall in this category are VilBert [7], VL-Bert [8] and Visual Bert [9].
Others, like LXMERT [10], Uniter [11], SimvLM [12], and most recently, OFA [13], use other transformer-based architectures. They all have in common that they are evaluated at least partly on one or more of the benchmark tasks provided by datasets like Visual Genome [14], MS-Coco [15], or Flickr30k [16]. Normally, the tasks are divided into two sub-categories which are briefly described here.
@@
This list is not exhaustive, but contains tasks most commonly referred to in Visual-Linguistic representation learning. The task with the highest similarity to the Perception-guided Crossmodal Entailment task we propose in this work is Visual Entailment [17]. In Visual Entailment, an image text combination functions as a premise/hypothesis pair, and the task is to classify whether the hypothesis is entailed in the premise, contradicts the premise, or is neutral with regard to the premise.
@@ 2.3 AI Alignment
The goal of AI alignment is to ensure that AI models behave in a way that overlaps with human preferences. Due to the ever increasing impact of LLMs and AI models in general, behavior such as halluzinations and social biases must be mitigated to prevent harmful effects [18]. To prevent such harmful responses, reinforcement learning from human feedback (RLHF) was originally introduced into the LLM deployment process. This technique relies on human-provided feedback, which serves as an alignment signal to prevent LLMs from returning harmful output [19] and has since seen continuous improvements.
@@ 2.4 Machine Learning on Eye Tracking Data
Historically, eye tracking data have been most common as a signal in psychology of (visual) perception, where it has been used to determine how visual perception works, both physiologically and psychologically [20]. In other lines of psychology-related work, eye tracking data are used to train statistical models to predict behavior in binary choice settings (cmp. [21] [22] [23]). In contrast to our work, the works described above use directly derived information from the eye tracking sequences (i.e. dwell time, revisits) [24] as input for their statistical models, while we focus on symbolic representations of multimodal AOIsthat we use as input for deep learning models.
@@
Most importantly, no previous work attempted to fuse the two research areas in order to derive individually aligned machine learning models.
@@ 3 Perception-guided Crossmodal Entailment
This section describes how the data for both the perception-guided ML model exploiting eye tracking data $X^{\text{\tiny ALIGN}}$, as well as the content-based transformer models trained solely on the stimulus data $X^{\text{\tiny STIM}}$ were obtained. For both models, identical prediction targets $Y^{\text{\tiny POV}}$ are used, according to the task of Perception-guided Crossmodal Entailment (PCE).
As a basis for the eye tracking study, we selected a set of multimodal documents from the Visual Genome data set [14]. Visual Genome (VG) consists of 108,249 images that are annotated on several layers, from regions to objects, represented in graph structures. We leverage the underlying scene graph to rank entities depicted in an image by their centrality degree. Based on this ranking, we chose regions that are connected by their contained entities and retrieved the associated caption. The captions of the regions then serve as a textual description of the picture. The region descriptions combined with the image constitute the multi-modal stimulus data $X^{\text{\tiny STIM}}$.
@@ 3.1 POV Task Specification
The eye tracking study consisted of 153 manually selected stimuli. After presenting $X^{\text{\tiny STIM}}$ to the study participants, we asked them the question “Are the central objects in the image mentioned in the caption?” resulting in their subjective evaluations $Y^{\text{\tiny POV}}$. We call this task Perception-guided Crossmodal Entailment (PCE) since it requires cross-modal reasoning and assesses individual human perception.
@@ 3.2 Eye tracking and Indvidual Assessment Recording
Overall, 109 participants in the age range of 19-61 years participated in the study, which took place in the eye tracking laboratory of the Media Studies department of our institution. Our institution’s ethics board approved the experiment as no harm was to be expected for the participants when evaluating image captions. The majority of the participants were female (F= 77, 70.64 %; M= 30, 27.52 %), the average age was Mage = 25.4 years. To ensure that the length of an eye tracking session was not too long to retain the participant’s attention, the participants were divided into three groups. A short instruction slide was shown followed by 50-53 blocks consisting of a text-image-stimulus and a following question slide. The stimulus and question slide blocks were shown in a randomized order. For the study, the experiment was programmed, recorded, and analyzed in the software iMotions (Version 9.1.5). The total duration for each session was approximately eight minutes (excluding the time for initial calibration).
@@ 3.3 Symbolic Fixation Sequence Extraction
From the eye tracking study we obtain participant-specific fixation sequences $X^{\text{\tiny ALIGN}}$ for each stimulus $X^{\text{\tiny STIM}}$. In addition to the information about the participant and stimulus, each fixation contains information about its location, its duration, dispersion, and the associated annotated region. For example, Table 4 lists how the sequence in Figure 4 is represented. All fixations are ordered by the time of their occurrence.
[…]
Nr. AOI Coordinates Duration
1 $wall_{vis}$ 865, 346 100.00
2 $wall_{text}$ 709, 29 158.32
20 $wall_{vis}$ 988, 431 250.00
Figure 3: Fixation sequence for Participant EWCX and Stimulus ID 2412873.
[…]
@@ 3.4 Data summary
In summary, the data we obtain consist of the created stimuli (images, annotated AOIs and corresponding captions) $X^{\text{\tiny STIM}}$ a participant-specific fixation sequence $X^{\text{\tiny ALIGN}}$ and an individual PCE assessment $Y^{\text{\tiny POV}}$ for each stimulus. Overall, the acquired data set contains 5,400 unique fixation sequences, where each sequence represents a perception pattern of a participant generated when parsing the stimulus. The fixation sequences contain a total of 148,100 identified fixations, on average 27.42 fixations per stimulus exposure. The distribution of the prediction targets $Y^{\text{\tiny POV}}$ is heavily skewed, with 67.4 % of the samples belonging to the Yes class, 20.1 % to the No class, and 12.5 % to the Unclear class. The data set (including raw fixation data, stimuli, caption texts, and annotated regions) and the code for the experiments is available on GitHub. [^1] https://github.com/siwer/Crossmodal-Entailment
@@ 4 POV Models and Experimental Design
To test our hypothesis that the stimuli evaluation process $X^{\text{\tiny STIM}}$ is reflected in the eye tracking data $X^{\text{\tiny ALIGN}}$ and can act as a feature to improve individual alignment $g^{\text{\tiny POV}}$, three models $f()$ were implemented with focus on different aspects of $X$.
The focus of the LSTM [25] model is on exploring the informational value of the fixation sequence of symbolic components of the multimodal stimulus as an alignment signal $x^{\text{\tiny ALIGN}_{\text{eye}}}$, without relying on content information $x^{\text{\tiny STIM}}$ on visual and textual entities in the stimuli. In contrast to this, the focus of the Transformer model is on content-based information in the form of pre-trained feature representations of the stimuli $x^{\text{\tiny STIM}}$. In this way, we can quantify the influence of both components of $x^{\text{\tiny POV}}$ independently of each other. To further test how individual preferences (without eye tracking information) impact performance, we also provide the participant ID as an alignment signal $x^{\text{\tiny ALIGN}_{\text{id}}}$ to the Transformer. Finally, as a combination of all, we train an Ensemble model that evaluates the predictions of all models together to see if the signals complement each other.
The task of all experimental setups is to predict the evaluation of each participant of the text-image combinations individually. This problem can be formulated as a classification problem with three possible outcomes for $y^{\text{\tiny POV}}$: 1) yes, the caption mentions the central objects in the image, 2) no, central objects in the image are not mentioned and 3) unclear.
4.1 Empirical Setup
The data are divided into training, evaluation and test sets in a random 80/10/10 split, with each split retaining the global class distribution from the data. All the models described in the following sections are optimized with AdamW [26] optimizer on a cross-entropy loss and trained for a maximum of 30 epochs. The ranges of hyperparameters that were tested to find the best performing Ensemble and best performing Perception-guided Transformer are: LR (1e-04, 5e-04, 1e-05, 5e-05), FF-Dim (32, 64, 128, 256), Embedding Dim (8, 16, 32), Batch size (16,64,128). All possible combinations were evaluated. We report accuracy and F1 score as evaluation metrics $g^{\text{\tiny POV}}$ to account for class imbalance in the data. Due to the sparsity of the unclear class, we tested all models in two different settings, one with all classes present (3 classes) and one with the unclear class excluded from the evaluation.